From mar@roe.ac.uk Tue May 20 10:46:36 2014 Date: Tue, 20 May 2014 10:46:22 +0100 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , Dave Morris , Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Alastair Edge , Andrew Lawrence , Bob Mann , Dante Minniti , James Dunlop , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Maria-Rosa Cioni , matt.jarvis@astro.ox.ac.uk, Mike Irwin , Nigel Metcalfe , Norman Gray , Phil Lucas , Richard McMahon at IOA , Simon Dye , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives progress meeting minutes, 9 May 2014 WFAU VDFS Science Archives progress meeting minutes, 9 May 2014 Present: MAR, RPB, ACD, ETWS, NJC, RGM Apologies: NCH, KTN, AL, RSC, MSH DoNM: 10am Friday 23 May 2014 in the Villas meeting room, MAR/ACD in the chair (room is booked). Actions discharged this time: ---------------------------- ACTION: MAR to email Monika Petr-Gotzens details on and examples of the ATLAS pointing offsets. Done. ACTION: ETWS to liaise with STV over accessing OSAProprietary (synced DB of metadata) through the new UI Done. ACTION: RGM to check on what is required re the Chilean U-band data. Done, and would be further discussed in the planned meetings next week. ACTION: MAR to apply P91 ATLAS QC. Done, the release is running. Actions partly discharged but continuing: ---------------------------------------- None. Actions carried forward from 25/4/2014 meeting: ---------------------------------------------- ACTION: ACD, RPB to formulate a policy for backups and for checking that data could be recovered from the backups. Continues. Notes on a proposed policy are available on the Twiki. No decision had yet been made on backing up the SVN repository. ACTION: RPB to put all prices of additional items associated with the backup policy on the Twiki pages. Continues. ACTION: RGM to investigate the size and content (how does it differ from DR9?) of DR10 to see if it's worth us hosting and whether it is necessary to keep all of the other releases. Continues. RGM had started a wiki page for WFAU SDSS strategy (http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan). ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this one still has low priority. ACTION: MAR to document issue with ATLAS unpaired sources and give examples of recovering matches. Continues. ACTION: ETWS to investigate modifying post CU4 ops to calculate and store illumination correction values. Continues. ACTION: ACD to arrange follow-up VVV brainstorming meeting. Continues Specific points and new actions: ------------------------------- Project management: The proposed meeting with Lockheed Martin and the UK Space Agency had been postponed. RGM noted that WFAU staff might have the opportunity to move to offices in the new Higgs building when completed. WFCAM, VISTA and VST updates: Nothing was raised. Comments and issues arising from CASU minutes: No minutes had been received since the previous meeting. Review of Deadlines: Those present went over the relevant entries on http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines ACTIION: MAR to look at the requirements and timescales for the next batch of ESO releases. Networking: Working. OSA data transferred up to March. WSA/VSA/OSA/GES Operations: The VVV objID update was still proceeding. The GPS nebulised data needed for DR10 have been transferred and ingested. Once checked, QC could be finished off and the release process started. ETWS noted that the recent WFCAM data that included a night of the satellite data held the satellite products in a sub-directory of the nightly directory. This results in the duplication of some files/filenames though it wasn't clear if the files had subtle differences and whether this was going to be the norm going forward. This would be raised when Mike Irwin visited next week. Hardware and Systems: MAR reported that a database on ramses11 and then ramses11 itself had gone awol. Re-attaching the database had cured the first issue but a power cycle had been "needed" to fix the second. Software: ACD noted that RSC had been working on modifying GES ingest code and that he'd been working on pre-ingest software. ETWS had modded syncDB to produce GES test DBs. Survey Data Release: RPB noted that VMCv20140438 (P91 v1.3) had been released. Non-survey Data Release: None since last time. Astrogrid deployment, VO & Data Analysis services: DM had been installing new TAP services in front of the earlier UKIDSS releases to test the ADQL parsing. Miscellaneous: VMC meeting at ROE 19-21 May 2014. -- Scanned by iCritical.